how do you simplify the ratio like 4kg/600g
Answer by venugopalramana(3286)

how do you simplify the ratio like 4kg/600g
IN CASE OF RATIOS ,BOTH NUMERATOR AND DENOMINATOR SHOULD BE IN SAME UNITS.
HERE NUMERATOR IS IN KG,WHERE AS D.R IS IN GMS.
SO CHANGE N.R. TO GMS.
4 KG=4KG.*1000GMS/1 KG=4000 GMS
HENCE RATIO IS = 4000/600=40/6=20/3